OpenStudy (anonymous):

(x + 3y = 2), (4x + 5y = 1) system of equations

OpenStudy (anonymous):

first, make the first equation x= x+3y=2 x=-3y+2 now plug that in for x in the second equation 4(-3y+2)+5y=1 -12y+8+5=1 -12y+13=1 -12y=-12 y=1 now plug y back into your x= equation x=-3(1)+2 x=-1 (-1,1)

OpenStudy (anonymous):

what about by elimination/addition?